# Accurate Power Spectrum Estimation toward Nyquist limit

### v1

@@ Abstract
The power spectrum, as a statistic in Fourier space, is commonly numerically calculated using the fast Fourier transform method to efficiently reduce the computational costs. To alleviate the systematic bias known as aliasing due to the insufficient sampling, the interlacing technique was proposed. We derive the analytical form of the shot noise under the interlacing technique, which enables the exact separation of the Poisson shot noise from the signal in this case. Thanks to the accurate shot noise subtraction, we demonstrate an enhancement in the accuracy of power spectrum estimation, and compare with the other widely used estimators in community. The good performance of our estimation allows an abatement in the computational cost by using low resolution and low order mass assignment scheme in the analysis for huge surveys and mocks.
@@ 1 Introduction
The power spectrum, usually denoted as $P(k)$ for three dimensional space, is an important statistical quantity in cosmological research. It is the Fourier transform pair of the two-point correlation function and reflects the magnitude of the matter density fluctuations at different spatial scales. The most accurate power spectrum estimation is the direct summation [9], but it is very time consuming as the computational cost is proportional to the number of grid points in ${\boldsymbol{k}}$ space as well as the number of galaxies. Thus, the power spectrum is usually calculated using the Fast Fourier Transform (FFT) method. By factorizing the calculation to even and odd series recursively, FFT reduces the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, with $N$ being the total number of grid points.
@@ 2.1 Power spectrum estimator
$W({\boldsymbol{k}})=\left[\frac{\sin(\pi k_{1}/k_{H})}{\pi k_{1}/k_{H}}\frac{\sin(\pi k_{2}/k_{H})}{\pi k_{2}/k_{H}}\frac{\sin(\pi k_{3}/k_{H})}{\pi k_{3}/k_{H}}\right]^{p}\ .$ (2.3)
For NGP, CIC, TSC and PSC, $p=1,2,3$ and $4$, respectively. $k_{H}=\frac{2\pi N_{g}}{L}$ represents the spatial frequency corresponding to the minimum grid point distance, and $N_{g}$ represents the number of grid points in one dimension. Following the derivation, conventions and expressions in [7], the resulting power spectrum is
$\langle|\delta^{f}({\boldsymbol{k}})|^{2}\rangle$ $=\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})+\frac{1}{N}\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}\ .$ (2.4)
@@ 2.2 No interlacing
As in [7], we denote the shot noise term as $\frac{1}{N}C({\boldsymbol{k}})\equiv\frac{1}{N}\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}$. From the identity
@@
the summation over infinite images of $W$ in one dimension can be derived. Let $q=2p$ and $x=\pi k/k_{H}$,
$\sum_{n}|W(k+k_{H}n)|^{2}$ $=\sum_{n}\left(\frac{\sin(\pi k/k_{H}+n\pi)}{\pi k/k_{H}+n\pi}\right)^{2p}$ (2.6)
@@
We denote the result as $S_{q}[\sin^{2}x]$ since it is a polynomial function of $\sin^{2}x$ up to order of $p-1$.
@@
$C({\boldsymbol{k}})=\left\{\begin{aligned} &1\ ,&\text{NGP}\ ,\\ &\prod_{i}\left[1-\frac{2}{3}\sin^{2}\left(\frac{\pi k_{i}}{k_{H}}\right)\right]\ ,&\text{CIC}\ ,\\ &\prod_{i}\left[1-\sin^{2}\left(\frac{\pi k_{i}}{k_{H}}\right)+\frac{2}{15}\sin^{4}\left(\frac{\pi k_{i}}{k_{H}}\right)\right]\ ,&\text{TSC}\ ,\\ &\prod_{i}\left[1-\frac{4}{3}\sin^{2}\left(\frac{\pi k_{i}}{k_{H}}\right)+\frac{2}{5}\sin^{4}\left(\frac{\pi k_{i}}{k_{H}}\right)-\frac{4}{315}\sin^{6}\left(\frac{\pi k_{i}}{k_{H}}\right)\right]\ ,&\text{PSC}\ .\\ \end{aligned}\right.$ (2.8)
@@
Neglecting the contamination from aliased images on the signal, i.e. assuming that $\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})\approx P({\boldsymbol{k}})|W({\boldsymbol{k}})|^{2}$, a straightforward estimator is
@@
Under a different approximation $\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})\approx P({\boldsymbol{k}})\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}=P({\boldsymbol{k}})C({\boldsymbol{k}})$, and thus $\langle|\delta^{f}({\boldsymbol{k}})|^{2}\rangle\approx[P({\boldsymbol{k}})+\frac{1}{N}]C({\boldsymbol{k}})$, the corresponding estimator is
@@
For this estimator, the window effect is corrected first and the $1/N$ shot noise is subtracted in the end. We caution the readers that the window effect here is $C({\boldsymbol{k}})$ instead of $W^{2}({\boldsymbol{k}})$. This correction can be performed on the field level, by dividing the effective window function $\sqrt{C({\boldsymbol{k}})}$ from the Fourier transformed field $\delta({\boldsymbol{k}})$. 0This is also the approach adopted in Nbodykit [^1] https://github.com/bccp/nbodykit, commit 376c9d7  [5] for the power spectrum estimation without interlacing technique.
@@ 2.3 Interlacing technique
Figure 1: Shot noise estimated from Poisson-distributed random samples. On the left side are the results for the CIC window function, and on the right side are the results for the TSC window function. In the top panels the scattered points represent the $\langle\delta^{f}(k)^{2}\rangle$ obtained from the random sample, the blue dashed lines follow Eq. (2.8) and the red lines follow Eq. (2.15). The middle panels show the results scaled to the Eq. (2.8) and to the Eq. (2.15) for the bottom panels. The analytical expressions well describe the shot noise with interlacing technique.
Hockney and Eastwood, [6] proposed the interlacing technique, which was lately applied to the $AP^{3}M$ N-body code [2]. This technique samples the signals twice with the second mesh having an overall half-cell offset $(H/2,H/2,H/2)$ relative to the primary mesh. The odd terms in the aliasing images will be eliminated by taking the arithmetic average of the two samplings. As such, the effects of aliasing can be reduced, leading to an improved estimation of the power spectrum. We denote the average of the two samplings as $\tilde{\delta}^{f}({\boldsymbol{k}})$, and mathematically,
@@
$=\sum_{\boldsymbol{n}}{\theta_{\boldsymbol{n}}{\delta}^{f}({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})}\ .$
@@
Again, ${\boldsymbol{n}}=0$ is the true signal term. In most cases and also true for cosmological matter power spectrum, the signal decreases with increasing scale, and thus the most dominant contribution is from $|{\boldsymbol{n}}|=1$. With the interlacing technique above, the aliasing effect from all odd pixels is eliminated. This approach smooths the sampling near the Nyquist frequency. At the same time, it should be noted that the shot noise is also affected by this smoothing effect.
The shot noise term with interlacing is $C^{{\operatorname{int}}}({\boldsymbol{k}})=\sum_{{\boldsymbol{n}}}\theta_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}$, in other words we are now interested in the summation of $\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}$ but over even $n_{x}+n_{y}+n_{z}$ terms. These terms can be split into two groups: all the $n_{i}$’s are even, or two of the three $n_{i}$’s are odd. Let $n=2m$, the summation over even terms in Eq. (2.6) is reduced to
@@
Here, $x=\frac{\pi k}{k_{H}}$. A similar derivation of the above relation is shown in [9] to obtain the relation between the residuals of the power spectrum estimator with and without interlacing. The summation over odd terms can be derived by letting $n=2m+1$,
@@
$C^{{\operatorname{int}}}({\boldsymbol{k}})=\prod_{i}C_{e}\left(\frac{\pi k_{i}}{k_{H}}\right)+\sum_{(i,j,m)\in A_{3}}C_{e}\left(\frac{\pi k_{i}}{k_{H}}\right)C_{o}\left(\frac{\pi k_{j}}{k_{H}}\right)C_{o}\left(\frac{\pi k_{m}}{k_{H}}\right)\ .$ (2.15)
@@
To exam the accuracy of the above shot noise formula, we calculate the power spectrum of a random sample with number density of $n_{s}=1.12\times 10^{-3}\text{ Mpc}^{-3}h^{3}$. We show the performance of the Eq. (2.15) in Figure 1, along with the result for Eq. (2.8) (also appeared in [7]). The analytical expressions describe the data points from the numerical calculation well, with the residual scattering around the expectation. The accuracy of the interlaced shot noise formula lines up with the accuracy of the case without interlacing.
@@
Under the same approximation, $\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})\approx P({\boldsymbol{k}})|W({\boldsymbol{k}})|^{2}$ in deriving the estimator Eq. (2.9), the new estimator for interlacing is
@@
The interlacing technique reduces aliasing both in the signal term and in the shot noise term. Under this consideration, we can directly discard the summation as $\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}\left[P({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})+\frac{1}{N}\right]\approx[P({\boldsymbol{k}})+\frac{1}{N}]|W({\boldsymbol{k}})|^{2}$, and thus the corresponding estimator is
@@
$\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})$ $\approx P({\boldsymbol{k}})|W({\boldsymbol{k}})|^{2}$ $\hat{P}({\boldsymbol{k}})=\frac{\langle|\delta^{f}({\boldsymbol{k}})|^{2}\rangle-\frac{1}{N}C({\boldsymbol{k}})}{|W({\boldsymbol{k}})|^{2}}$, Eq. (2.9) Exact shot noise subtraction without interlacing.
$\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})$ $\approx P({\boldsymbol{k}})\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}$ $\hat{P}({\boldsymbol{k}})=\frac{\langle|\delta^{f}({\boldsymbol{k}})|^{2}\rangle}{C({\boldsymbol{k}})}-\frac{1}{N}$, Eq. (2.10) Window effect $\sqrt{C({\boldsymbol{k}})}$ can be deconvolved at field level.
$\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})$ $\approx P({\boldsymbol{k}})|W({\boldsymbol{k}})|^{2}$ $\hat{P}({\boldsymbol{k}})=\frac{\langle|\tilde{\delta}^{f}({\boldsymbol{k}})|^{2}\rangle-\frac{1}{N}C^{{\operatorname{int}}}({\boldsymbol{k}})}{|W({\boldsymbol{k}})|^{2}}$, Eq. (2.16) Exact shot noise subtraction for interlacing.
$\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}\left[P({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})+\frac{1}{N}\right]$ $\approx\left[P({\boldsymbol{k}})+\frac{1}{N}\right]|W({\boldsymbol{k}})|^{2}$ $\hat{P}({\boldsymbol{k}})=\frac{\langle|\tilde{\delta}^{f}({\boldsymbol{k}})|^{2}\rangle}{|W({\boldsymbol{k}})|^{2}}-\frac{1}{N}$, Eq. (2.17) Window effect $|W({\boldsymbol{k}})|$ can be deconvolved at field level.

### v2

@@ Abstract
The power spectrum, as a statistic in Fourier space, is commonly numerically calculated using the fast Fourier transform method to efficiently reduce the computational costs. To alleviate the systematic bias known as aliasing due to the insufficient sampling, the interlacing technique was proposed. We derive the analytical form of the shot noise under the interlacing technique, which enables the exact separation of the Poisson shot noise from the signal in this case. Thanks to the accurate shot noise subtraction, we demonstrate an enhancement in the accuracy of power spectrum estimation. For the three dark matter samples with number density ranging from $10^{-2}\text{ Mpc}^{-3}h^{3}$ to $10^{-4}\text{ Mpc}^{-3}h^{3}$ analyzed on a coarse mesh of size $3\,h^{-1}{\rm Mpc}$ with Cloud-in-Cell mass assignment scheme, the systematic bias is well under control up to the Nyquist frequency. On the contrary, the bias induced by the estimator that ignore the aliasing on the shot noise typically exceeds the statistical uncertainty on the frequency beyond 0.85 times the Nyquist frequency. The good performance of our estimation allows an abatement in the computational cost by using low resolution and low order mass assignment scheme in the analysis for huge surveys and mocks.
@@ 1 Introduction
The power spectrum, usually denoted as $P(k)$ for three dimensional space, is an important statistical quantity in cosmological research. It is the Fourier transform counterpart of the two-point correlation function and reflects the magnitude of the matter density fluctuations at different spatial scales. The most accurate power spectrum estimation is the direct summation [9], but it is very time consuming as the computational cost is proportional to the number of grid points in ${\boldsymbol{k}}$ space times the number of galaxies. Thus, the power spectrum is usually calculated using the Fast Fourier Transform (FFT) method. By factorizing the calculation to even and odd series recursively, FFT reduces the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, with $N$ being the total number of grid points.
@@ 2.1 Power spectrum estimator
$W({\boldsymbol{k}})=\left[\frac{\sin(\pi k_{1}/2k_{N})}{\pi k_{1}/2k_{N}}\frac{\sin(\pi k_{2}/2k_{N})}{\pi k_{2}/2k_{N}}\frac{\sin(\pi k_{3}/2k_{N})}{\pi k_{3}/2k_{N}}\right]^{p}\ .$ (2.3)
For NGP, CIC, TSC and PSC, $p=1,2,3$ and $4$, respectively. $k_{N}=\frac{\pi N_{g}}{L}$ represents the Nyquist frequency, and $N_{g}$ represents the number of grid points in one dimension. Following the derivation, conventions and expressions in [7], the resulting power spectrum is
$\langle|\delta^{f}({\boldsymbol{k}})|^{2}\rangle$ $=\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})+\frac{1}{N}\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}\ .$ (2.4)
@@ 2.2 No interlacing
As in [7], we denote the shot noise term as $\frac{1}{N}C({\boldsymbol{k}})\equiv\frac{1}{N}\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}$. From the identity
@@
the summation over infinite images of $W$ in one dimension can be derived. Let $q=2p$ and $x=\pi k/2k_{N}$,
$\sum_{n}|W(k+2k_{N}n)|^{2}$ $=\sum_{n}\left(\frac{\sin(\pi k/2k_{N}+n\pi)}{\pi k/2k_{N}+n\pi}\right)^{2p}$ (2.6)
@@
We denote the result as $S_{q}[\sin^{2}x]$ since it is a polynomial function of $\sin^{2}x$ up to order of $p-1$,
@@
$C({\boldsymbol{k}})=\left\{\begin{aligned} &1\ ,&\text{NGP}\ ,\\ &\prod_{i}\left[1-\frac{2}{3}\sin^{2}\left(\frac{\pi k_{i}}{2k_{N}}\right)\right]\ ,&\text{CIC}\ ,\\ &\prod_{i}\left[1-\sin^{2}\left(\frac{\pi k_{i}}{2k_{N}}\right)+\frac{2}{15}\sin^{4}\left(\frac{\pi k_{i}}{2k_{N}}\right)\right]\ ,&\text{TSC}\ ,\\ &\prod_{i}\left[1-\frac{4}{3}\sin^{2}\left(\frac{\pi k_{i}}{2k_{N}}\right)+\frac{2}{5}\sin^{4}\left(\frac{\pi k_{i}}{2k_{N}}\right)-\frac{4}{315}\sin^{6}\left(\frac{\pi k_{i}}{2k_{N}}\right)\right]\ ,&\text{PSC}\ .\\ \end{aligned}\right.$ (2.8)
@@
Neglecting the contamination from aliased images on the signal, i.e. assuming that $\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})\approx P({\boldsymbol{k}})|W({\boldsymbol{k}})|^{2}$, a straightforward estimator is
@@
Under a different approximation $\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})\approx P({\boldsymbol{k}})\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}=P({\boldsymbol{k}})C({\boldsymbol{k}})$, and thus $\langle|\delta^{f}({\boldsymbol{k}})|^{2}\rangle\approx[P({\boldsymbol{k}})+\frac{1}{N}]C({\boldsymbol{k}})$, the corresponding estimator is
@@
For this estimator, the window effect is corrected first and the $1/N$ shot noise is subtracted in the end. We caution the readers that the window effect here is $C({\boldsymbol{k}})$ instead of $W^{2}({\boldsymbol{k}})$. This correction can be performed on the field level, by dividing the effective window function $\sqrt{C({\boldsymbol{k}})}$ from the Fourier transformed field $\delta({\boldsymbol{k}})$. This is also the approach adopted in Nbodykit [^1] https://github.com/bccp/nbodykit, commit 376c9d7  [5] for the power spectrum estimation without interlacing technique.
@@ 2.3 Interlacing technique
Figure 1: The comparison of the analytical shot noise expressions with the measurement from Poisson-distributed random samples. On the left side are the results for the CIC window function, and on the right side are the results for the TSC window function. In the top panels the data points represent the $\langle\delta^{f}(k)^{2}\rangle/W^{2}(k)$ obtained from the random sample, the blue dashed lines follow Eq. (2.8) and the red lines follow Eq. (2.15). The middle panels show the results scaled to the Eq. (2.8) and to the Eq. (2.15) for the bottom panels. The analytical expressions well describe the shot noise with interlacing technique.
[6] proposed the interlacing technique, which was lately applied to the $AP^{3}M$ N-body code [2]. This technique samples the signals twice with the second mesh having an overall half-cell offset $(H/2,H/2,H/2)$ relative to the primary mesh. The odd terms in the aliasing images will be eliminated by taking the arithmetic average of the two samplings. As such, the effects of aliasing can be reduced, leading to an improved estimation of the power spectrum. We denote the average of the two samplings as $\tilde{\delta}^{f}({\boldsymbol{k}})$, and mathematically,
@@
$=\sum_{\boldsymbol{n}}{\theta_{\boldsymbol{n}}{\delta}^{f}({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})}\ .$
@@
Again, ${\boldsymbol{n}}=0$ is the true signal term. In most cases and also true for cosmological matter power spectrum, the signal decreases with increasing frequency, and thus the most dominant contribution is from $|{\boldsymbol{n}}|=1$. With the interlacing technique above, the aliasing effect from all odd pixels is eliminated. This approach smooths the sampling near the Nyquist frequency. At the same time, it should be noted that the shot noise is also affected by this smoothing effect.
The shot noise term with interlacing is $C^{{\operatorname{int}}}({\boldsymbol{k}})=\sum_{{\boldsymbol{n}}}\theta_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}$. In other words, we are now interested in the summation of $\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}$ but over even $n_{x}+n_{y}+n_{z}$ terms. These terms can be split into two groups: all the $n_{i}$’s are even, or two of the three $n_{i}$’s are odd. Let $n=2m$, the summation over even terms in Eq. (2.6) is reduced to
@@
Here, $x=\frac{\pi k}{2k_{N}}$. A similar derivation of the above relation is shown in [9] to obtain the relation between the residuals of the power spectrum estimator with and without interlacing. The summation over odd terms can be derived by letting $n=2m+1$,
@@
$C^{{\operatorname{int}}}({\boldsymbol{k}})=\prod_{i}C_{e}\left(\frac{\pi k_{i}}{2k_{N}}\right)+\sum_{(i,j,m)\in A_{3}}C_{e}\left(\frac{\pi k_{i}}{2k_{N}}\right)C_{o}\left(\frac{\pi k_{j}}{2k_{N}}\right)C_{o}\left(\frac{\pi k_{m}}{2k_{N}}\right)\ .$ (2.15)
@@
To exam the accuracy of the above shot noise formula, we calculate the power spectrum of a random sample with number density of $n_{s}=1.12\times 10^{-3}\text{ Mpc}^{-3}h^{3}$. We show the performance of the Eq. (2.15) in Figure 1, along with the result for Eq. (2.8) (also appeared in [7]). The results are presented in the form of $C(k)/W^{2}(k)$ to line up with the second term, i.e. the assumed shot noise contribution, in estimator Eq. (2.9). The analytical expressions describe the data points from the numerical calculation well, with the residual scattering around the expectation. The accuracy of the interlaced shot noise formula lines up with the accuracy of the case without interlacing.
@@
Under the same approximation, $\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})\approx P({\boldsymbol{k}})|W({\boldsymbol{k}})|^{2}$ in deriving the estimator Eq. (2.9), the new estimator for interlacing is
@@
The interlacing technique reduces aliasing both in the signal term and in the shot noise term. Under this consideration, we can directly discard the summation as $\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}\left[P({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})+\frac{1}{N}\right]\approx[P({\boldsymbol{k}})+\frac{1}{N}]|W({\boldsymbol{k}})|^{2}$, and thus the corresponding estimator is
@@
$\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})$ $\approx P({\boldsymbol{k}})|W({\boldsymbol{k}})|^{2}$ $\hat{P}({\boldsymbol{k}})=\frac{\langle|\delta^{f}({\boldsymbol{k}})|^{2}\rangle-\frac{1}{N}C({\boldsymbol{k}})}{|W({\boldsymbol{k}})|^{2}}$, Eq. (2.9) Exact shot noise subtraction without interlacing.
$\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})$ $\approx P({\boldsymbol{k}})\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}$ $\hat{P}({\boldsymbol{k}})=\frac{\langle|\delta^{f}({\boldsymbol{k}})|^{2}\rangle}{C({\boldsymbol{k}})}-\frac{1}{N}$, Eq. (2.10) Window effect $\sqrt{C({\boldsymbol{k}})}$ can be deconvolved at field level.
$\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}P({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})$ $\approx P({\boldsymbol{k}})|W({\boldsymbol{k}})|^{2}$ $\hat{P}({\boldsymbol{k}})=\frac{\langle|\tilde{\delta}^{f}({\boldsymbol{k}})|^{2}\rangle-\frac{1}{N}C^{{\operatorname{int}}}({\boldsymbol{k}})}{|W({\boldsymbol{k}})|^{2}}$, Eq. (2.16) Exact shot noise subtraction for interlacing.
$\sum_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}\left[P({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})+\frac{1}{N}\right]$ $\approx\left[P({\boldsymbol{k}})+\frac{1}{N}\right]|W({\boldsymbol{k}})|^{2}$ $\hat{P}({\boldsymbol{k}})=\frac{\langle|\tilde{\delta}^{f}({\boldsymbol{k}})|^{2}\rangle}{|W({\boldsymbol{k}})|^{2}}-\frac{1}{N}$, Eq. (2.17) Window effect $|W({\boldsymbol{k}})|$ can be deconvolved at field level.